OpenStudy (jellybot23):

Any help? solve 10=3^x+5 -2

OpenStudy (whpalmer4):

That equation does not look reasonable. Are you sure you have copied it correctly?

OpenStudy (jellybot23):

yes, it is 10=3^x+5 -2 The -2 is separate from the exponent

OpenStudy (whpalmer4):

you can't write it like that if you mean something other than \[10 = 3^x + 5 -2\]

OpenStudy (whpalmer4):

Is it supposed to be \[10 = 3^{x+5}-2\]

OpenStudy (jellybot23):

yes! the 2nd one :)

OpenStudy (whpalmer4):

If that's what you intended, you need to put parentheses around the exponent, like this: 10 = 3^(x+5) - 2 remember the order of precedence: PEMDAS parentheses exponentiation multiplication division addition subtraction so 10 = 3^x + 5 - 2 first operation done is 3^x, then you add 5, then you subtract 2

OpenStudy (jellybot23):

oh okay. Then I meant 10=3^(x+5) -2 Sorry about that

OpenStudy (whpalmer4):

usually, if you copy some equation where there are exponents or anything that relies on positioning, like a complicated fraction such as \[\frac{a+b}{c+d}\]you need to use some parentheses to make it clear (or the equation editor button). a lot of people might right that fraction as a+b/c + d but that really means \[a + \frac{b}{c} +d \]as they have written...

OpenStudy (whpalmer4):

but now that we have that out of the way, on to your problem :-) \[10 = 3^{x+5} - 2\]what do you think a good first step might be?

OpenStudy (jellybot23):

Okay, I will remember that :) And, well I thought you had to do \[\log_{3}10 = x+5\] first

OpenStudy (whpalmer4):

wait, what happened to the -2?

OpenStudy (jellybot23):

I was going to subtract two after I got my answer. I'm not very good at logs :( They always confuse me

OpenStudy (whpalmer4):

Can't do that. But even if you could, why wouldn't you get rid of that 2 first by adding 2 to both sides and simplifying the field of battle?

OpenStudy (jellybot23):

Oh true! so 12=3^(x+5)

OpenStudy (jellybot23):

Then do I do \[\log_{3}12=x+5 \]

OpenStudy (whpalmer4):

yes. keep going...

OpenStudy (jellybot23):

I plugged it into my calculator and got approximately -2.738

OpenStudy (whpalmer4):

And if you plug that value into the right hand side of the original equation, do you get a number very close to 10?

OpenStudy (jellybot23):

yes! :)

OpenStudy (whpalmer4):

Excellent! So, I agree that logs can be confusing, but if you think of them as the reverse of exponentiation, that helps me, at least.

OpenStudy (jellybot23):

Awesome! Thank you so so much :D

OpenStudy (whpalmer4):

you're welcome!